# Share Your Secrets for Privacy! Confidential Forecasting with Vertical Federated Learning

### v1

@@ Abstract
Vertical federated learning (VFL) is a promising area for time series forecasting in industrial applications, such as predictive maintenance and machine control. Critical challenges to address in manufacturing include data privacy and over-fitting on small and noisy datasets during both training and inference. Additionally, to increase industry adaptability, such forecasting models must scale well with the number of parties while ensuring strong convergence and low-tuning complexity. We address those challenges and propose “Secret-shared Time Series Forecasting with VFL" (STV), a novel framework that exhibits the following key features: i) a privacy-preserving algorithm for forecasting with SARIMAX and autoregressive trees on vertically-partitioned data; ii) serverless forecasting using secret sharing and multi-party computation; iii) novel $N$-party algorithms for matrix multiplication and inverse operations for direct parameter optimization, giving strong convergence with minimal hyperparameter tuning complexity. We conduct evaluations on six representative datasets from public and industry-specific contexts. Our results demonstrate that STV’s forecasting accuracy is comparable to those of centralized approaches. They also show that our direct optimization can outperform centralized methods, which include state-of-the-art diffusion models and long-short-term memory, by 23.81% on forecasting accuracy. We also conduct a scalability analysis by examining the communication costs of direct and iterative optimization to navigate the choice between the two. Code and appendix are available: https://github.com/adis98/STV.
[…]
2484
@@ 1 Introduction
Time series forecasting with vertically-partitioned data is critical in manufacturing applications like continuous operations, predictive maintenance, and machine control [22, 36]. For example, let us consider the scenario in Figure 1 with two suppliers (factories) and their customer. The suppliers produce components that go into the manufacturing of a device possessed by the customer. They collect sensory data during production while the customer owns post-production performance data, serving as the outputs or labels. Forecasting the outputs from sensor values would allow the suppliers to make pre-shipment corrections, saving valuable time and helping to calibrate equipment in a complex, high-volume production factory. Each factory may possess a different set of sensors, i.e., different input feature sets as they all produce different components. Training a model on the combined feature sets could boost the predictive capability of the forecasting model as there may be associations between the feature sets of different clients. This calls for a collaborative forecasting framework that jointly trains a model using all feature sets to predict the performance. However, the main challenge with training such a joint model is data privacy. Specifically, confidentiality agreements between parties or legal regulations that prevent the sharing of sensitive performance and sensor data.
Existing framework. To address these privacy concerns, federated learning (FL) has been emerging as a promising solution [18]. In FL, training follows a model-to-data approach without data leaving the party’s premises. In Vertical federated learning (VFL), each participant owns a different feature set pertaining to the same sample ID [44]. This is the case with our manufacturing example, since each manufacturer has a different set of sensors, each one corresponding to a different set of attributes in the finished product.
@@
Figure 1: Problem scenario—forecasting device performance needs inputs from multiple parties, all of whom want to protect the confidentiality of their data.
Challenges. Despite its relevance, time-series forecasting with VFL has received limited attention [43]. This underscores the critical need for further exploration and research in this domain, especially considering three challenges that manufacturing scenarios introduce.
First, deep neural networks (DNNs) tend to overfit on manufacturing datasets since they can be affected by slow collection and noisy measurements, leading to small datasets [21, 11, 46]. Moreover, their complex parameter interactions limit their interpretability, which is needed for understanding the reasoning behind predictions to optimize production quality and downtime [41, 40].
Second, VFL is predominated by approaches utilizing a hierarchical split-learning architecture. This consists of several bottom models at the clients and a single top model held by a server [38, 16, 7]. The bottom models compute forward activations of their party’s sensitive features. The top model is held by a server that combines the intermediate activations to produce the output. Any party requiring forecasts would need the server to first produce the output, i.e., act as a middleman. This strong dependence on a single party is infeasible for several reasons. The server is a single point of failure which is catastrophic in case of data breaches. Moreover, it requires strong trust from the other parties as they are heavily reliant upon it for forecasting outputs. This is difficult to negotiate in manufacturing environments involving business competitors.
Third, while analytical/direct optimization methods like the normal equation [4] reach globally optimal solutions without requiring any hyperparameter tuning, they do not scale well to large problem instances, unlike iterative methods such as gradient descent. However, industries require both scalable and convenient solutions as they encounter diverse scenarios.
Contributions. To address these challenges, we develop a novel framework, Secret-Shared Time Series Forecasting with Vertical Federated Learning (STV), with the following contributions.
1. VFL forecasting framework. STV enables forecasting for VFL using Secret sharing (SS) [33] and secure multi-party computation (SMPC) [9, 24]. These do not involve privacy-performance trade-offs and offer cryptographic privacy guarantees. We propose STVL for linear models such as SARIMAX [19, 29, 13], and STVT for autoregressive trees (ARTs) [27].
2. Private, serverless inference. By adopting SS and SMPC, all computations and outputs are performed in a decentralized fashion by the parties involved. Thus, the outputs and intermediate data exist as distributed shares across all parties. STV ensures that the party requiring inference obtains the final predictions first-hand by collecting these shares. Due to the dispersion of shares, responsibility of output generation is distributed across all the involved parties, promoting trust through mutual dependence.
3. Adaptable optimization with Least Squares. STVL uses a two-step approach with least squares (LS). This lends adaptability, as LS can be optimized using both iterative and direct methods. This requires protocols for $N$-party matrix multiplications and inverses on secretly shared data, a key novelty in our work. While iterative approaches like gradient descent scale better, direct optimization methods do not require hyperparameter tuning and guarantee global convergence. Therefore, we offer both options for tackling diverse forecasting scenarios.
@@
We thoroughly evaluate STV on multiple fronts. First, we compare the forecasting accuracy of STVL and STVT with centralized state-of-the-art forecasters based on diffusion models [2], Long Short Term Memory (LSTM), and SARIMAX with Maximum Likelihood Estimation (MLE) [13]. Second, we compare the communication costs of iterative and direct optimization of linear forecasters under different scaling scenarios, highlighting their trade-offs. We use a wide range of datasets: five public datasets (Air quality, flight passengers, SML 2010, PV Power, Rossman Sales) and one real dataset from semiconductor manufacturing for measuring chip overlays from noisy alignment sensor data.
@@ 2 Background
2.1 Time series forecasting
STV uses a popular linear forecaster, SARIMAX (Seasonal Auto-Regressive Integrated Moving Average with eXogenous variables) [19, 29], that generalizes other autoregresive forecasters [13, 28]. It models the series to forecast as a linear function composed of autoregressive, moving averages, and exogenous variables along with their seasonal counterparts [29].
Autoregressive terms capture the influence of earlier series outputs on future values. The moving average terms capture the influence of earlier errors/residuals. Finally, exogenous features serve as additional external features that may aid with forecasting. The seasonal counterparts model periodic patterns in the series. For example, output $Y$ at time $t$ can be represented using a polynomial $H$, containing residuals $\epsilon(t-i)$, past values $Y(t-i)$, and exogenous features $X_{j}(t)$ using a linear equation as follows:
The coefficients ($\alpha,\beta,\gamma$) can be estimated using least-squares (LS) [15, 26, 25, 37] or Maximum Likelihood Estimation (MLE) [13]. However, implementing MLE with multiparty computation and secret sharing is limited to likelihood functions like the exponential or multivariate normal distributions [35, 23], since computations with SMPC are limited to a small number of supported mathematical operations. However, least-squares optimization is still possible by extending basic SMPC protocols such as scalar addition and multiplication (see Section 2.2). First, the datasets, $(X,Y)$, are transformed into time-lagged design matrices, ($\phi_{X},\phi_{Y}$), to represent Equation 1:
In least-squares optimization, $A$ can be optimized using a two-step regression approach [37, 26, 15, 25]. First, the residuals are estimated by modeling using only autoregressive (AR) and exogenous terms. Then, all the coefficients in $A$ are jointly optimized by setting the residuals to the estimates.
@@ 2.1 Time series forecasting
Normal equation:
$\hat{A}=(({\hat{\phi_{X}}})^{T}\hat{\phi_{X})}^{-1}(({\hat{\phi_{X}}})^{T}{\phi_{Y}})`$ (5)
Then, the residuals, $\varepsilon$, are then estimated as follows:
@@
These residual estimates are then re-substituted in Equation 2 to refine $A$ via a second optimization step.
While the above example is applicable to linear forecasters, the idea of autoregression can also be utilized for tree-based models like XGBoost [6]. This leads to the formulation of autoregressive trees (ARTs) [27], which use previous output values as decision nodes in a gradient-boosted regression tree. Therefore, extending VFL methods for XGBoost [42, 8, 10] to ARTs requires transforming the datasets into design matrices using a polynomial like Equation 1. While ARTs do not use residual terms, the tree-based architecture enables modeling forecasts in a non-linear fashion.
2.2 SMPC
@@
Parties cannot infer others’ data from their shares alone; however, the value can be recovered by combining all shares. By analogy, we distribute local features and outputs as shares to preserve their privacy. All parties then jointly utilize decentralized training protocols on the secretly shared data to obtain a local model. Inference/forecasting is done by distributing features into secret shares and then computing the prediction as a distributed share across all parties.
Training and forecasting on secretly shared features requires primitives for performing mathematical operations on distributed shares, which we explain as follows.
@@
Multiplication (using Beaver’s triples) [3, 42]: Consider $Z=X*Y$, where * denotes element-wise multiplication, and $X$ and $Y$, are secretly shared. The coordinator first generates three numbers $a,b,c$ such that $c=a*b$. These are then secretly shared, i.e., $C_{k}$, receives $\langle a\rangle^{k}$, $\langle b\rangle^{k}$, and $\langle c\rangle^{k}$. $C_{k}$ computes $\langle e\rangle^{k}$ = $\langle X\rangle^{k}$ $\langle a\rangle^{k}$ and $\langle f\rangle^{k}$ = $\langle Y\rangle^{k}$ $\langle b\rangle^{k}$, and sends it to $C_{1}$. $C_{1}$ then aggregates these shares to recover $e$ and $f$ and broadcasts them to all parties. $C_{1}$ then computes $\langle Z\rangle^{1}$ = $e*f$+$f*$$\langle a\rangle^{1}$+ $e*$$\langle b\rangle^{1}$+$\langle c\rangle^{1}$, and the others calculate $\langle Z\rangle^{k}$ = $f*$$\langle a\rangle^{k}$+$e*$$\langle b\rangle^{k}$+$\langle c\rangle^{k}$. It is easy to see that aggregation of the individual shares gives the product $Z$.
Despite knowing $e$ and $f$, $X$ and $Y$ are hidden since the parties do not know $a,b,$ and $c$. Similar to addition, the individual shares, $\langle Z\rangle^{k}$, do not reveal anything about the local share values, i.e., $\langle X\rangle^{k}$, $\langle Y\rangle^{k}$, $\langle a\rangle^{k}$, $\langle b\rangle^{k}$, and $\langle c\rangle^{k}$.
Additional primitives such as division, argmax, and sigmoid can also be computed using secret sharing, the details of which are provided in Fang et al. [10] and Xie et al. [42].
@@ 3 STV Framework
(a) Polynomial generation through pre-preocessing
@@
Figure 3: Decentralized training and inference in STV
@@
Output:Trained model distributed across $K$ parties or final inference result on party $C_{j}$
@@
7:  $\langle\phi_{X}\rangle$,$\langle\phi_{Y}\rangle$ = TransformData ($H$){ Equation 2}
@@
12:     else if $O=``direct"$ then
@@
Here we introduce the adversarial model and problem statement, followed by STV’s design and implementation for ART, STVT, and SARIMAX forecasters, STVL. For both models, secret sharing and SMPC are used to protect the privacy of input features.
Adversarial Models. We assume that all parties are honest-but-curious/semi-honest [16, 44], i.e., they adhere to protocol but try to infer others’ private data using their own local data and whatever is communicated to them. Also, it is assumed that parties do not collude. This is a standard assumption in VFL as all parties are incentivised to collaborate due to their mutual dependence on one another for training and inference [44, 16]. In addition, we also assume that communication between parties is encrypted.
Problem Statement. We assume a setup with $K$ parties, $C_{1}$ to $C_{K}$, grouped into two types: $active$ and $passive$. The active party, $C_{1}$, owns the outputs of the training data for the time series, $Y(t)$, for timestep $t$. It also possesses exogenous features, $X_{1}(t)$. The passive parties only have exogenous features, $X_{i},\forall i\in[2,K]$. The common samples between parties are assumed to be already identified using privacy-preserving entity alignment approaches [16, 32]. Our goal is to forecast future values using exogenous and autoregressive features without sharing them with others in plaintext.
We further assume that there is a coordinator, a trusted third party that oversees the training process and is responsible for generating randomness, such as Beaver’s triples for element-wise multiplication with SMPC [3, 10]. The coordinator cannot access private data and intermediate results, so it does not pose a privacy threat, as mentioned by Xie et al. [42].

### v2

@@ Abstract
Vertical federated learning (VFL) is a promising area for time series forecasting in many applications, such as healthcare and manufacturing. Critical challenges to address include data privacy and over-fitting on small and noisy datasets during both training and inference. Additionally, such forecasting models must scale well with the number of parties while ensuring strong convergence and low-tuning complexity. We address these challenges and propose “Secret-shared Time Series Forecasting with VFL” (STV), a novel framework with the following key features: i) a privacy-preserving algorithm for forecasting with SARIMAX and autoregressive trees on vertically-partitioned data; ii) decentralised forecasting using secret sharing and multi-party computation; and iii) novel $N$-party algorithms for matrix multiplication and inverse operations for exact parameter optimization, giving strong convergence with minimal tuning complexity. We evaluate on six representative datasets from public and industry-specific contexts. Results demonstrate that STV’s forecasting accuracy is comparable to those of centralized approaches. Our exact optimization outperforms centralized methods, including state-of-the-art diffusion models and long-short-term memory, by 23.81% on forecasting accuracy. We also evaluate scalability by examining the communication costs of exact and iterative optimization to navigate the choice between the two. STV’s code and supplementary material is available online: https://github.com/adis98/STV.
Keywords:
Vertical Federated Learning Time Series ForecastingMultiparty Computation
@@ 1 Introduction
Vertically-partitioned or feature-partitioned time series data are prevalent in many areas, such as healthcare, manufacturing, and finance [46, 40, 30]. For example, consider the scenario in Figure 1, where a mental health facility and a cardiac centre possess distinct features, such as stress levels and heart rate corresponding to a common patient. Each party can forecast patient risk levels using just their own feature sets, but it is likely that combining stress levels and heart rate could lead to better predictive precision due to the correlations between the two [10]. The straightforward way to use both sources is to first centralize the two feature sets at a server and then train a model. However, privacy restrictions, such as patient confidentiality agreements or laws (e.g. GDPR [33]), prevent data sharing between organizations. Such a scenario naturally extends to other domains, such as joint predictive maintenance in manufacturing[23, 39], or stock predictions in finance[30].
Federated learning (FL) [19] is a promising research direction to address the privacy concern. Its training paradigm follows a model-to-data approach where data never leaves the premises of the source. Within FL, Vertical federated learning (VFL), considers cases where each participant owns different features pertaining to the same sample ID or timestamp [46], which corresponds to our problem scenario. However, time-series forecasting with VFL has received limited attention [45, 40] and existing works overlook the following challenges.
First, VFL methods typically employ deep learning, which suffers when data is scarce due to overfitting. Data scarcity is a real issue—healthcare and manufacturing environments can be affected by slow collection and noisy measurements, leading to small datasets [22, 47, 26]. Second, VFL methods predominantly use a split-learning architecture, consisting of several bottom models at the clients and a single top model held by a server [42, 17, 7]. Final predictions are assumed to be generated by the top model at the server. However, since each party may have an interest in maintaining control or ownership over the final output, deciding which party should assume the server role can lead to conflicts. Moreover, centralizing predictions at a single party would be catastrophic in case of a data breach. Third, the models are generally trained iteratively—e.g., via gradient descent—needing extensive hyperparameter tuning. In contrast, analytical/exact methods can reach globally optimal solutions without any tuning, but may incur higher compute complexity. Therefore, we need a way to flexibly incorporate both approaches to handle diverse scenarios.
@@
Figure 1: Problem scenario—forecasting risk levels needs corresponding input features from multiple parties, all of whom want to protect the confidentiality of their data.
To address the challenges, we present a novel framework, Secret-Shared Time Series Forecasting with Vertical Federated Learning (STV), with the following contributions. 1. VFL forecasting framework—built using secret sharing (SS) [36] and multi-party computation (MPC) [9, 25] for cryptographic privacy guarantees. We propose STVL for linear models such as SARIMAX [20, 32, 14], and STVT for autoregressive trees (ARTs) [29]. 2. Decentralized inference—all outputs and intermediate data remain distributed across parties as shares when using MPC. Hence, responsibility of output generation is also distributed across all the involved parties, promoting trust through mutual dependence. 3. Adaptable optimization with Least Squares—using a two-step approach that can flexibly be incorporated into both iterative and exact methods. This approach requires algorithms for $N$-party matrix multiplications and inverses on secret shares, a key novelty in our work.
@@
Our evaluation compares the forecasting accuracy of STVL and STVT with centralized state-of-the-art forecasters based on diffusion models [2], Long Short Term Memory (LSTM), and SARIMAX with Maximum Likelihood Estimation (MLE) [14]. We also compare the communication costs of iterative and exact optimization of linear forecasters under different scaling scenarios, highlighting their trade-offs. We use a wide range of datasets: five public datasets (Air quality, flight passengers, SML 2010, PV Power, Rossman Sales) and one industrial semiconductor dataset for measuring chip overlays from alignment sensor data.
@@ 2 Background
2.1 Time series forecasting
STV uses a popular linear forecaster, SARIMAX (Seasonal AutoRegressive Integrated Moving Average with eXogenous variables) [20, 32], that generalizes other autoregresive forecasters, such as ARMA, ARIMA, and ARIMAX [14, 31]. It models the outputs as a linear function composed of autoregressive, moving averages, and exogenous variables along with their seasonal counterparts [32].
Autoregressive terms capture the influence of historical values on future predictions and moving average terms capture the influence of historical errors/residuals. Exogenous features serve as additional auxiliary features that aid in forecasting. The seasonal terms capture periodic patterns in the series. Mathematically, the output $Y$ at time $t$ can be modelled using a linear polynomial $H$, containing historical residuals $\epsilon(t-i)$, historical observations $Y(t-i)$, and exogenous features $X(t)$ as follows:
The coefficients ($\alpha,\beta,\gamma$) are generally estimated using least-squares [16, 28, 27, 41] or Maximum Likelihood Estimation (MLE) [14]. However, implementing MLE with multiparty computation and secret sharing is limited to likelihood functions like the exponential or multivariate normal distributions [38, 24], since computations with MPC are limited by a small number of supported mathematical operations. However, least-squares optimization is still possible by extending basic MPC protocols such as scalar addition and multiplication (see Section 2.2). First, we transform the datasets, $(X,Y)$, into time-lagged design matrices ($\phi_{X},\phi_{Y}$) representing eq.(1):
@@ 2.1 Time series forecasting
With least-squares, $A$ can be optimized using a two-step regression approach [41, 28, 16, 27]. First, the residuals are estimated by modeling using only autoregressive (AR) and exogenous terms. Then, all the coefficients in $A$ are jointly optimized by setting the residuals to the estimates.
@@
Following this initial estimate, the residuals, $\varepsilon$, are then obtained as follows:
@@
These residual estimates are then re-substituted in eq.(2) to refine $\hat{A}$ via a second optimization step. Although we describe the two-step optimization procedure in the context of linear forecasters, the idea of autoregression can also be extended to tree-based models like XGBoost [6]. Autoregressive trees (ARTs) [29] build on this premise and use historical outputs/observations as decision nodes in a gradient-boosted regression tree. Therefore, extending XGBoost-based VFL methods [44, 8, 11] to ARTs only requires transforming the datasets into lagged design matrices using a polynomial like eq.(1). This architecture enables modelling forecasts in a non-linear fashion.
2.2 MPC
@@
Parties cannot infer others’ data from their shares alone, as shares are completely random. However, the value can be recovered by combining all shares. We can extend this idea to machine learning by distributing feature vectors and outputs as shares to preserve their privacy. All parties then jointly utilize decentralized training protocols on the secretly shared data to obtain a local model. Inference/forecasting is then done by distributing features into secret shares and then computing the prediction as a distributed share across all parties. Training and forecasting on secretly shared features requires primitives for performing mathematical operations on distributed shares, which we explain as follows.
@@
Multiplication (using Beaver’s triples) [3, 44]: Consider $Z=X*Y$, where * denotes element-wise multiplication, and $X$ and $Y$, are secretly shared. The coordinator first generates three numbers $a,b,c$ such that $c=a*b$. These are then secretly shared, i.e., $C_{k}$, receives $\langle a\rangle^{k}$, $\langle b\rangle^{k}$, and $\langle c\rangle^{k}$. $C_{k}$ computes $\langle e\rangle^{k}$ = $\langle X\rangle^{k}$ $\langle a\rangle^{k}$ and $\langle f\rangle^{k}$ = $\langle Y\rangle^{k}$ $\langle b\rangle^{k}$, and sends it to $C_{1}$. $C_{1}$ then aggregates these shares to recover $e$ and $f$ and broadcasts them to all parties. $C_{1}$ then computes $\langle Z\rangle^{1}$ = $e*f$+$f*$$\langle a\rangle^{1}$+ $e*$$\langle b\rangle^{1}$+$\langle c\rangle^{1}$, and the others calculate $\langle Z\rangle^{k}$ = $f*$$\langle a\rangle^{k}$+$e*$$\langle b\rangle^{k}$+$\langle c\rangle^{k}$. It is easy to see that aggregation of the individual shares gives the product $Z$. Despite knowing $e$ and $f$, $X$ and $Y$ are hidden since the parties do not know $a,b,$ and $c$. Moreover, the values of $X$ and $Y$ are also hidden from the coordinator, who is only responsible for the generation of $a,b$, and $c$, and does not hold any features or outputs. Similar to addition, the individual shares, $\langle Z\rangle^{k}$, do not reveal anything about the local share values, i.e., $\langle X\rangle^{k}$, $\langle Y\rangle^{k}$, $\langle a\rangle^{k}$, $\langle b\rangle^{k}$, and $\langle c\rangle^{k}$. Additional primitives for division and argmax can also be computed using MPC, as detailed in Fang et al. [11] and Xie et al. [44].
3 Related Work
Table 2: Comparison of related works in VFL.
Method
Time Series
Serverless inference
Dual optimization
N-party ($\geq 2$)
  Yan et al. [45]
Han et al.[15]
Xie et al.[44]
Shi et al.[37]
STV (this work)
As shown in Table 2, we compare VFL methods on their applicability to time series forecasting, potential for achieving inference privacy, adaptability to alternative optimization choices, and generalizability to $N$-parties.
Time-series forecasting. Earlier, we mentioned that industries require easy-to-understand and low-complexity models to avoid overfitting on small datasets. Hence we focus on linear/logistic regression (LR) [17, 15, 37], and tree-based models [44, 8]. Yan et al. [45] modifies the split learning architecture using Gated Recurrent Units (GRUs) with a shared upper model for predictions. However, it depends on a single server to produce forecasts, which is a bottleneck and can lead to trust issues among the involved parties.
Decentralized inference. We consider schemes adopting SS as potential candidates for our inference requirements, as they are straightforward to integrate into a decentralized approach like ours [44, 37, 15]. Shi et al. [37] and Han et al. [15] train linear models while Xie et al. [44] implement XGBoost. Homomorphic encryption methods, such as Hardy et al. [17], require the predictions to be decrypted at the server, violating the requirement.
Optimization. When it comes to optimization techniques, all selected works, except for Han et al. [15], employ solely iterative approaches. Notably, Han et al. [15] offer iterative and matrix-based methods for exact optimization using eq.(4). However, they only provide 2-party protocols for matrix multiplications and inverses, which we extend to the $N$-party case.
@@ 4 STV Framework
(a) Polynomial generation through pre-processing
@@
Figure 3: Training and inference in STV
@@
Output:Trained model distributed across $K$ parties or predicted forecast at party $C_{j}$
@@
7:  $\langle\phi_{X}\rangle$,$\langle\phi_{Y}\rangle$ = TransformData ($H$){ (2)}
@@
12:     else if $O=``exact"$ then
@@
Here we introduce the adversarial model and problem statement, followed by STV’s design and implementation for trees (STVT), and SARIMAX (STVL). Both models employ secret sharing and MPC to protect the privacy of features.
Adversarial Models. We assume that all parties are honest-but-curious/semi-honest [17, 46], i.e., they adhere to protocol but try to infer others’ private data using their own local data and whatever is communicated to them. Also, it is assumed that parties do not collude. This is a standard assumption in VFL as all parties are incentivised to collaborate due to their mutual dependence on one another for training and inference [46, 17]. In addition, we also assume that communication between parties is encrypted to prevent snooping.
Problem Statement. We assume a setup with $K$ parties, $C_{1}$ to $C_{K}$, grouped into two types: $active$ and $passive$. Without loss of generality, we denote the active party as $C_{1}$, which holds the ground truth outputs of the training data for the time series, $Y(t)$, for a timestep $t$. It also possesses exogenous features, $X_{1}(t)$. The passive parties only have exogenous features, $X_{i},\forall i\in[2,K]$. The common samples between parties are assumed to be already identified using privacy-preserving entity alignment approaches [17, 35]. Our goal is to forecast future values using exogenous and autoregressive features without sharing them with others in plaintext. We further assume that there is a coordinator, a trusted third party that oversees the training process and is responsible for generating randomness, such as Beaver’s triples for element-wise multiplication with MPC [3, 11]. However, this coordinator cannot access private data and intermediate results, so it does not pose a privacy threat, as mentioned in Xie et al. [44].